The following information pertains to the inventory of Parvin Company during Year 2:

Jan. 1 Beginning Inventory 400 units @ $ 30
Apr. 1 Purchased 2,150 units @ $ 35
Oct. 1 Purchased 600 units @ $ 38


During Year 2, Parvin sold 2,900 units of inventory at $90 per unit and incurred $47,000 of operating expenses. Parvin currently uses the FIFO method but is considering a change to LIFO. All transactions are cash transactions. Assume a 30 percent income tax rate. Parvin started the period with cash of $86,000, inventory of $12,000, common stock of $61,000, and retained earnings of $37,000.

Exercise 5-6A Part c

c. Determine the amount of income tax that Parvin would pay using each cost flow method.
  

income tax paid using fico
incometax paid using lico
Answer #1
c
Sales 261000 =2900*90
Less: Cost of goods sold FIFO 100550 =(400*30)+(2150*35)+(350*38)
Gross margin 160450
Less: Operating expenses 47000
Income before income tax 113450
Income tax paid using FIFO 34035 =113450*30%
Sales 261000
Less: Cost of goods sold LIFO 102550 =(600*38)+(2150*35)+(150*30)
Gross margin 158450
Less: Operating expenses 47000
Income before income tax 111450
Income tax paid using LIFO 33435 =111450*30%